# MACE-OFF: Short Range Transferable Machine Learning Force Fields for Organic Molecules

### v3

@@ III.3.2 Lattice enthalphies
Next, we assess the ability of the MACE-OFF23 force fields to describe the stability of molecular crystals. We computed the enthalpies of sublimation of a range of 23 representative small molecular crystals [78] following the protocol of Ref. [79].
Figure 6 compares the predicted sublimation enthalpies with the experimentally measured ones. This task is often used to test various DFT functionals. Since the ωB97M-D3(BJ) functional used to parameterize MACE-OFF23 does not have a periodic implementation, an estimate of the highest achievable accuracy is not available. The figure shows that the three MACE models are able to capture trends and improve significantly compared to the ANI-2x model. The large MACE-OFF23 model achieves a mean absolute error of just 1.7 kcal/mol, which is comparable to the errors of several different dispersion-corrected density functionals for a fraction of the computational cost [78].
@@
Figure 6: Sublimation enthalphies of molecular crystals. Comparison between predicted sublimation enthalphies of the MACE and ANI models and experiment.
We have also compared the relaxed unit cell vectors of the MACE-OFF23 models with the experimentally measured values and found that they have a relative error as low as 5% with the MACE-OFF23(L) model as shown in the supporting Table 9.
III.4 Water structure and dynamics
A key requirement for bio-organic force fields is an accurate description of water. Several empirical water models have been developed for this purpose, the most common for all-atom biomolecular simulations being TIP3P. Figure 7 shows that MACE-OFF23(M) correctly predicts the radial distribution function (RDF) of water, performing comparably with both TIP3P and the highly accurate MB-pol water model, whose RDF is indistinguishable from experiment [80]. MACE-OFF23(S) gives a qualitatively similar RDF, with one noteable shift in the position of the second peak in the $g_{OO}(r)$ peak from 4.5 Å to 4 Å. This performance is consistent with the results discussed in more detail in section III.6. These interactions are modelled in a purely local way by MACE, having been trained on small water clusters containing up to 50 water molecules. This result further emphasizes the ability of the model to generalize to bulk simulations, even though the training data only contained non-periodic clusters. In contrast, ANI-2X shows a significant over-structuring of the RDF, likely due to the lack of dispersion correction in the underlying quantum mechanical method used to label the training set.
We further calculated the temperature dependence of the density using NPT molecular dynamics simulations, as shown in Figure 9. This is a highly sensitive test for the accuracy of intermolecular interactions [65]. Both the small and medium models over-predict the density to a small extent, which likely originates from the underlying DFT functional or the lack of periodic training data.
Next, we characterize the dynamical properties of water by computing the vibrational density of states, including the IR and Raman spectra selection rules, using MACE models of dipoles and polarizabilities [73]. Given the significant impact of quantum nuclear motion on the dynamics of water [71], we incorporate quantum nuclear effects using the PIGS approach [71] (discussed in Section III.3.1), which has a classical cost and has recently been shown to describe quantum nuclear effects in water accurately [81]. As shown in Figure 7, after incorporating quantum nuclear effects, the MACE-OFF23(S) predictions agree with the experiments across the entire frequency range. We note the presence of an overall 2-3 % blue shift with respect to the experiments, consistent with the spectra for paracetamol form II (see Section III.3.1). MACE-OFF23(S) even qualitatively captures subtle features of the spectra, including the 0 – 1000 $\text{cm}^{-1}$ hydrogen bonding fingerprints in the anisotropic Raman spectrum, and the bimodal nature of the isotropic Raman stretching band arising from the presence of hydrogen-bonded defects in room temperature water.
Figure 7: Water structure and dynamics. Top: IR, isotropic Raman and anisotropic Raman spectra of water at ambient temperature and density using the MACE-OFF23(S) model for the potential energy surface and a single MACE model of water’s dipole and polarizability [73]. The black curve represents the experiment [82]. The dashed blue curve is obtained from standard (classical) MD, while the solid curve incorporates quantum effects using the PIGS method [71]. Bottom: Radial distribution function of MACE, ANI and TIP3P models. MB-pol traces are reproduced from Ref. [80].
III.5 Condensed phase properties of molecular liquids
Small molecule empirical force fields are fitted to reproduce experimental condensed phase properties including densities and heats of vaporization. For machine learning force fields, this task represents a greater challenge, since the potential is typically trained only using QM reference data for isolated molecules and molecular dimers. From these ab initio data, the models learn the long-range interactions necessary to reproduce bulk properties without ever fitting to them.
@@ III.5 Condensed phase properties of molecular liquids
To investigate the performance of the MACE-OFF23 models in the condensed phase, we selected a benchmark set of 109 molecules from Ref. [83], containing a representative set of functional groups relevant to medicinal chemistry and biology. MACE-OFF23(M) has a larger receptive field and is therefore better suited for reproducing experimental observables most affected by intermolecular interactions. In general, we found that the small model gave a reasonable description in the NVT ensemble, while the medium model is also able to accurately model liquids in NPT simulations, as detailed below. For the liquid boxes containing approximately 600 atoms, MACE-OFF23(S) achieved a throughput of $2.1\times 10^{6}$ steps/day via OpenMM on a single 80GB NVIDIA A100 GPU, whilst the medium model achieved a throughput of $1.1\times 10^{6}$ steps/day. Meanwhile MACE-OFF23(L) achieved a throughput of $2.8\times 10^{5}$ steps/day on the same system, highlighting the significant overhead compared to the other two models. See section III.7 for further discussion of the computational performance.
III.5.1 Organic liquids
First, we investigate the predictions of molecular liquid densities under ambient conditions using the medium MACE model shown on the top panel of Figure 8. MACE-OFF23(M) achieves a MAE of 0.09 g/cm3, indicating that it can generalize from the intermolecular interactions between molecular dimers in the training set to the condensed phase, retaining a good correlation with experiment. In contrast, the ANI-2x potential has significantly higher errors, with MAE and RMSE errors twice as large as MACE-OFF23(M).
@@ III.5.1 Organic liquids
Interestingly, we found that MACE-OFF23(S) systematically overpredicts the experimental densities (SI Table 6) with a MAE of 0.23 g/cm3. This is likely the result of the slightly higher energy error and the shorter cutoff radius. In particular, the contribution of the intermolecular interactions to the overall energy is much smaller than the intramolecular interactions making the fractional error in them much larger. This is demonstrated in SI Table 4, where the test errors are decomposed into inter- and intramolecular contributions. We find that the medium and large models have significantly lower intermolecular force errors. Whilst we see MACE-OFF23(L) as a replacement for semi-empirical QM due to its relative expense compared to the small and medium models, we nonetheless benchmarked the large model on a small selection of condensed phase systems. We saw comparable performance to the medium model, and in particular observed that compounds that had a high prediction error under the medium model had a similar error under the large model. This indicates that this error is driven primarily by the dataset rather than by the size of the model, as previously hypothesized.
We further investigate the performance of MACE-OFF23(M) on heats of vaporization. The predictions correlate strongly with the experimental data); however, there is a systematic offset of approximately 2 kcal/mol. We hypothesized that this may result from an underprediction of the total intermolecular interactions originating from two sources: those that appear in molecular dynamics simulations outside the effective cutoff of the MACE model (compare classical simulations, which typically employ long-range corrections to the total energy [84]), and those that the model has not learned from the limited set of intermolecular interactions in the training set.
To investigate this further, we reran the calculations with MACE-OFF24(M) (see SI), which employs an extended cutoff of 6 Å and has also been trained on additional nonbonded data. We observe a modest decrease in MAE from 2.18 to 1.75 kcal/mol, as well as a small decrease in the accuracy of liquid density predictions. This further reinforces the hypothesis that the poorly predicted densities are a result of under-representation of the functional groups, since the additional training sets do little to increase sampling of these problematic functional groups.
Interestingly, for MACE-OFF23(S), this offset disappears (see SI Table 7), however this is likely due to the overpredicted densities, resulting in a cancellation of errors through overprediction of the magnitude of intermolecular interactions.
Nevertheless, these data indicate that, given a reasonable receptive field and a sufficiently expressive model, it is possible to fit a local force field that captures the interactions required to recover experimental properties.
III.6 Biomolecular simulations
@@
III.6.1 Extended cutoff MACE-OFF models
One key observation that emerged from validation of the MACE-OFF23(S) and MACE-OFF23(M) models was an overestimation of the density of liquid water by approximately 20% (Figure 9). Since there is extensive coverage of water clusters in the training set, we hypothesized that this resulted from a missing long-range contribution beyond the receptive field. To investigate this hypothesis, we fitted an additional model, MACE-OFF23(M)b, which has an increased 6 Å layerwise cutoff compared to 5 Å. The larger receptive field leads to a predicted water density within 2% of the experimental value at room temperature.
We also investigated the effect of including additional training data, made up of the Solvated Pubchem and Amino Acid Ligand Pairs datasets from SPICE v2 [63]. Given the direct applicability of these training data to biomolecular simulations, in the following section we test an additional model, MACE-OFF24(M), which includes the additional training data. Given the encouraging performance of MACE-OFF23(M)b, this model also employs the extended 6 Å cutoff.
Figure 9: Water density as a function of temperature for the extended cutoff MACE-OFF models
III.6.2 Ala3 Free Energy Surface
@@
Figure 10 shows the free energy surface of Ala3 in water, modelled by AMBER14/TIP3P and MACE-OFF24(M), respectively. Well-tempered metadynamics was used to enhance sampling along the central $\phi$ and $\psi$ backbone angles. The AMBER14 simulation identifies four local minima, corresponding to the anti-parallel $\beta$-sheet ($\phi<-120^{\circ},\psi>120^{\circ}$), a right-handed $\alpha$-helix ($\phi=-60^{\circ},\psi>-60^{\circ}$), the corresponding left-handed $\alpha$-helix ($\phi=60^{\circ},\psi>60^{\circ}$), and a polyproline II (PPII)-type structure ($\phi=-60^{\circ},\psi>120^{\circ}$). Comparing the dihedral angle distribution with NMR J-coupling constants, the AMBER14 simulation is generally in good agreement, with a slight overpopulation of both $\alpha$-helical structures and underpopulation of the PPII mesostate [85]. We therefore use this to enable visual comparison with a well-tuned classical forcefield.
The corresponding MACE simulations show a very similar free energy surface to AMBER14/TIP3P, with the same four local minima identified. Whilst the relative depths of the α-helical conformations are in agreement, AMBER predicts the free energy of the anti-parallel β-sheet to be 1 kcal/mol higher than the PPII minimum, compared to 0.1 kcal/mol for MACE.
We also compare the unbiased conformational ensemble predicted by MACE directly to experiment by calculating 3J coupling constants from the Ramachandran distribution. Computed values are shown in Table 3 in addition to literature values for AMBER14SB, ANI-2x and experiment [85, 86, 87]. MACE-OFF24(M) accurately predicts coupling constants, and is comparable with AMBER14SB. In agreement with our previous findings, both methods significantly outperform ANI-2x, highlighting the difficulties in computing accurate dynamics in the solution phase with MLPs.
3J(HN,Hα) 3J(HN,Cβ) 3J(HN,C)
MACE-OFF24(M) 5.70 1.64 1.37
ANI-2x 4.5 1.2 1.4
AMBER14SB 6.07 1.87 1.14
EXP 5.68 2.39 1.29
Table 3: Summary of 3J couplings computed from dihedral distributions. AMBER14SB and Experimental values taken from Ref. [85]. ANI-2x data taken from Ref. [86].
@@
III.6.3 Folding dynamics of Ala15
Having confirmed that MACE-OFF24(M) is capable of recovering the free energy surface of a small peptide, we next investigated the folding of the longer helical peptide Ala15. Because this is a significantly larger system size than the dipeptide configurations on which the model was trained, this represents a nontrivial test of the extrapolation capability of the potential, including its ability to capture complex hydrogen bonding interactions required to stabilize the secondary structure. Such simulations are believed to be particularly difficult with purely local models such as those in the MACE-OFF family.
As an initial test, the fully extended Ala15 structure was simulated in vacuo. Figure 11 shows the assignment of the secondary structure during the simulation. The peptide folded within 200 ps, adopting initially a “wavy” intermediate structure, before first folding into a 310 helix. The secondary structure oscillates between the α-helix and the 310 helix for the remainder of the simulation, with the α-helix being the predominant motif. These results are in agreement with other works [9, 39], in which polyalanine peptides are observed to fold through a “wavy intermediate” and oscillate between 310 and α-helices. The oscillatory behaviour has also been observed experimentally [88].
To better understand the performance of MACE-OFF24(M), we compared the predictions to those of MACE-OFF23(M) and MACE-OFF23(M)b. Through ablation studies, we found that both the inclusion of additional training configurations and the increased radial cutoff were required to observe the oscillatory behavior, whilst all other models overstabilize the α-helix, and do not predict the 310 structure. We hypothesize that additional sampling of the nonbonded interactions present in the expanded training set, as well as a larger receptive field of 12 Å allows the model to learn more subtle hydrogen bonding interactions required to interconvert between the two helical motifs.
We also measured the secondary structure of the peptide as the temperature was increased. Starting from the folded structure at 300K, the temperature was increased in steps from 300K to 900K over the course of 3 ns (Figure 12). The folded structure remains stable up to around 500 K, after which point the helix unfolds and the structure is primarily a random coil. This is an underprediction compared to the experimental unfolding temperature of 725 K, and suggests that there are additional subtle stabilising interactions that are not fully captured by the model [89].
Figure 11: Ala15 folding dynamics. Secondary structure assignment during 500 ps trajectory generated with MACE-OFF24(M)FF24M. The structure proceeds via a “wavy” intermediate and oscillates between the α- and 310 helices. Secondary structure assignment was performed with the STRIDE algorithm [90].
@@ III.6.3 Folding dynamics of Ala15
III.6.4 Protein simulation in explicit solvent
Finally, we investigated the simulation of a fully solvated protein with MACE-OFF. As a test case, we chose Crambin, a 42 residue threonin storage protein that contains four charged residues. This test goes significantly beyond what can typically be expected from a local force field trained on neutral species only.
Using the MACE-OFF24(M) model, we first confirmed that the force field was indeed capable of simulating a large biomolecular system containing charged residues. In particular, it has recently been shown that the FENNIX potential fails to simulate a solvated 28 residue peptide due to unphysical proton transfers involving charged residues [37]. We confirmed that the root mean square fluctuations of the protein backbone relative to the minimized structure remained below 1 Å for the duration of a 1 ns simulation, no bond breaking occurred and the secondary structure motifs remained intact (see SI Figure 17).
We then investigated the ability of MACE-OFF24(M) to capture the key vibrational modes of the system. To this end, we prepared a simulation box containing crambin in the high-hydration state, that is, fully solvated in explicit water, for a total system size of approximately 18,000 atoms. The power spectrum was calculated from the final 125 ps of a 1.6 ns simulation recorded at 2 fs resolution, to match the simulation length and sampling frequency employed in Ref [41]. Experimentally, this system has been well studied, and key THz region vibrational modes have been assigned [91]. This system has also been investigated with the AMBER classical force field and several machine learned potentials, including GEMS and, more recently, SO3LR [40, 39, 9]. All three MLPs produce spectra in qualitative agreement with each other. The water peaks at 1640 cm-1 and 3200-3600 cm-1 are reproduced reasonably accurately, especially considering that no nuclear quantum effects are taken into account (Figure 13). Furthermore, in the terahertz region, two peaks are identified by both GEMS, SO3LR and MACE-OFF24(M). The most obvious of these occurs at 60 cm-1 and corresponds to localised internal side chain fluctuations, while a peak at 220 cm-1 is due to specific configurations of water at the protein surface.
We note that, whilst a 125 ps trajectory is sufficient to qualitatively identify the spectral features, a much longer simulation of at least several nanoseconds would be required to fully converge the higher frequency THz region of the spectrum and enable full quantitative comparison with experiment, as demonstrated for the GEMS model in Ref. [39]. We note that the power spectrum computed on the final 1.3 ns simulation (discarding the initial 300 ps to equilibration) does not yield a significantly different spectrum to that of the 125 ps simulation used in Figure 13, indicating that significantly more sampling would be required to resolve the THz lineshapes sufficiently to enable experimental comparison.

### v4

@@ III.3.1 Lattice enthalphies
We assess the ability of the MACE-OFF23(M) model to describe the stability of molecular crystals. We computed the enthalpies of sublimation of a range of 23 representative small molecular crystals [73] following the protocol of Ref. [74] (Section S4).
[…]
Additional analyses of the relaxed unit cell vectors using the MACE-OFF23(L) model, and the powder Raman spectrum of crystalline paracetamol are presented in Sections S5 and S8.
@@
Figure 4: Sublimation enthalphies of molecular crystals. Comparison between predicted sublimation enthalphies of the MACE-OFF23(M) and ANI models and experiment.
III.3.2 Condensed phase properties of organic liquids
@@
Small molecule empirical force fields are fitted to reproduce experimental condensed phase properties including densities and heats of vaporization. As such, errors on small molecular systems are expected to be well within chemical accuracy [77, 21]. For machine learning force fields, this task represents a greater challenge, since the potential is typically trained only using QM reference data for isolated molecules and molecular dimers. From these ab initio data, the models must learn the long-range interactions necessary to reproduce bulk properties without ever fitting to them.
To investigate the performance of the MACE-OFF23 models in the condensed phase, we selected a benchmark set of 109 molecules from Ref. [78], containing a representative set of functional groups relevant to medicinal chemistry and biology. For the liquid boxes containing approximately 600 atoms, MACE-OFF23(S) achieved a throughput of $2.1\times 10^{6}$ steps/day via OpenMM on a single 80GB NVIDIA A100 GPU, whilst the medium model achieved a throughput of $1.1\times 10^{6}$ steps/day. Meanwhile MACE-OFF23(L) achieved a throughput of $2.8\times 10^{5}$ steps/day on the same system, highlighting the significant overhead compared to the other two models. See section III.5 for further discussion of the computational performance.
First, we investigate the predictions of molecular liquid densities under ambient conditions using the medium MACE model shown on the top panel of Figure 5. MACE-OFF23(M) achieves a MAE of 0.09 g/cm3, indicating that it can generalize from the intermolecular interactions between molecular dimers in the training set to the condensed phase, retaining a good correlation with experiment. In contrast, the ANI-2x potential has significantly higher errors, with MAE and RMSE errors twice as large as MACE-OFF23(M) (Section S6).
@@
We further investigated the performance of MACE-OFF23(M) on heats of vaporization. The predictions correlate strongly with the experimental data; however, there is a systematic offset of approximately 2 kcal/mol (Figure 5). We hypothesized that this may result from an under-prediction of the total intermolecular interactions originating from two sources: those that appear in molecular dynamics simulations outside the effective cutoff of the MACE model (compare classical simulations, which typically employ long-range corrections to the total energy [79]), and those that the model has not learned from the limited set of intermolecular interactions in the training set.
The effect of the choice of MACE-OFF model on condensed phase properties is further investigated in Section S6. In general, we observe that the small model systematically over-predicts the experimental density (MAE 0.23 g/cm3), which is consistent with its larger intermolecular force errors (Table S1). We did not observe any significant improvements in accuracy, over the medium model, using either the large model or a model trained with an extended cutoff of 6 Å. This further reinforces the hypothesis that the poorly predicted densities are a result of under-representation of the functional groups in the training set. Nevertheless, these data indicate that, given a reasonable receptive field and a sufficiently expressive model, it is possible to fit a local force field that captures the interactions required to recover experimental properties. Further improvements in accuracy are possible without resorting to fitting to experiment by increasing further the size of the training dataset and the explicit addition of long-range Coulomb interactions into the model, which we leave for future work.
III.4 Biomolecular simulations
@@
III.4.1 Water structure and dynamics
A key requirement for bio-organic force fields is an accurate description of water. Several empirical water models have been developed for this purpose, the most common for all-atom biomolecular simulations being TIP3P. Figure 6 shows that MACE-OFF23(M) correctly predicts the radial distribution function (RDF) of water, performing comparably with both TIP3P and the highly accurate MB-pol water model, whose RDF is indistinguishable from experiment [80]. These interactions are modelled in a purely local way by MACE, having been trained on small water clusters containing up to 50 water molecules. This result further emphasizes the ability of the model to generalize to bulk simulations, even though the training data only contained non-periodic clusters. In contrast, ANI-2X shows a significant over-structuring of the RDF, however this is likely driven primarily by the lack of dispersion correction in the underlying quantum mechanical method used to label the training set. Additional tests of the vibrational density of states of liquid water are presented in Section S8 and show good agreement with experiment.
Figure 6: Water structure. Radial distribution functions of MACE, ANI and TIP3P models. MB-pol traces are reproduced from Ref. [80].
We further calculated the temperature dependence of the water density using NPT molecular dynamics simulations, as shown in Figure 7. This is a highly sensitive test of the accuracy of intermolecular interactions [65]. The MACE-OFF23(M) model over-estimates the density of liquid water by approximately 20%. Since there is extensive coverage of water clusters in the training set, we hypothesized that this results from a missing long-range contribution beyond the receptive field. To investigate this hypothesis, we fitted an additional model, which has an increased 6 Å layerwise cutoff compared to 5 Å. The larger receptive field leads to a predicted water density within 2% of the experimental value at room temperature, and within 5% across the full temperature range (Figure 7). We note that similar errors are reported using many dispersion-corrected DFT exchange-correlation functionals [81]. At this point in the model development, the second version of SPICE had been released, which included Solvated Pubchem and Amino Acid Ligand Pairs datasets [63]. Given the direct applicability of these training data and the importance of modelling water to biomolecular simulations, the remainder of this section uses the MACE-OFF24(M) model, which includes the additional training data and extended 6 Å cutoff (Table 2). To confirm no loss of performance on incorporation of the additional training data, we confirmed that the test set errors are in agreement with MACE-OFF23(M) (Section S2.1).
Figure 7: Water density as a function of temperature for MACE-OFF23(M), and the extended cutoff MACE-OFF24(M) model.
III.4.2 Ala3 Free Energy Surface
@@
Figure 8 shows the free energy surface of Ala3 in water, modelled by AMBER14SB/TIP3P and MACE-OFF24(M), respectively. Well-tempered metadynamics was used to enhance sampling along the central $\phi$ and $\psi$ backbone angles (Section S7.1). The AMBER14SB simulation identifies four local minima, corresponding to the anti-parallel $\beta$-sheet ($\phi<-120^{\circ},\psi>120^{\circ}$), a right-handed $\alpha$-helix ($\phi=-60^{\circ},\psi>-60^{\circ}$), the corresponding left-handed $\alpha$-helix ($\phi=60^{\circ},\psi>60^{\circ}$), and a polyproline II (PPII)-type structure ($\phi=-60^{\circ},\psi>120^{\circ}$). Comparing the dihedral angle distribution with NMR J-coupling constants, the AMBER14SB simulation is generally in good agreement, with a slight overpopulation of both $\alpha$-helical structures and underpopulation of the PPII mesostate [82]. We therefore use this to enable visual comparison with a well-tuned classical forcefield.
[…]
We also compare the unbiased conformational ensemble from 20 ns of NPT dynamics sampled with MACE-OFF24(M) directly to experiment by calculating 3J coupling constants from the Ramachandran distribution. Computed values are shown in Table 3 in addition to literature values for AMBER14SB, ANI-2x and experiment [82, 83, 84]. MACE-OFF24(M) accurately predicts coupling constants, and is comparable with AMBER14SB. In agreement with our previous findings, both methods significantly outperform ANI-2x, highlighting the difficulties in computing accurate dynamics in the solution phase with MLPs.
[…]
MACE-OFF24(M) 5.70(0.43) 1.64(0.18) 1.37(0.30)
ANI-2x 7.5 1.2 1.8
AMBER14SB 6.07 1.87 1.13
EXP 5.68(0.11) 2.39(0.09) 1.13(0.08)
Table 3: Summary of 3J couplings computed from dihedral distributions. AMBER14SB and experimental values are taken from Ref. [82], and ANI-2x data from Ref. [83]. Standard errors in the mean from block averaging are shown in parentheses for MACE-OFF, and experimental uncertainties are reproduced from Ref. [82]
@@
III.4.3 Folding dynamics of Ala15
Having confirmed that MACE-OFF24(M) is capable of recovering a good approximation to the free energy surface of a small peptide, we next investigated the folding of the longer helical peptide Ala15. Because this is a significantly larger system size than the dipeptide configurations on which the model was trained, this represents a nontrivial test of the extrapolation capability of the potential, including its ability to capture complex hydrogen bonding interactions required to stabilize the secondary structure. Such simulations are believed to be particularly difficult with purely local models such as those in the MACE-OFF family.
[…]
We also measured the secondary structure of the peptide as the temperature was increased. Starting from the folded structure at 300 K, the temperature was increased in steps from 300 K to 900 K over the course of 3 ns (Figure 10). The folded structure remains stable up to around 500 K, after which point the helix unfolds and the structure is primarily a random coil.
[…]
@@ III.4.3 Folding dynamics of Ala15
III.4.4 Protein simulation in explicit solvent
Finally, we investigated the simulation of a fully solvated protein with MACE-OFF. As a test case, we chose Crambin, a 42-residue threonin storage protein that contains four charged residues. This test goes significantly beyond what can typically be expected from a local force field trained on neutral species only.
Using the MACE-OFF24(M) model, we first confirmed that the force field was indeed capable of simulating a large biomolecular system containing charged residues. In particular, it has recently been shown that the FENNIX potential fails to simulate a solvated 28 residue peptide due to unphysical proton transfers involving charged residues [37]. We confirmed that the root mean square fluctuations of the protein backbone relative to the minimized structure remained below 1 Å for the duration of a 1 ns simulation, no bond breaking occurred and the secondary structure motifs remained intact (Figure S2). We obtained a throughput of $3\times 10^{5}$ steps per day on a single NVIDIA A100 80GB GPU.
We then investigated the ability of MACE-OFF24(M) to capture the key vibrational modes of the system. To this end, we prepared a simulation box containing crambin in the high-hydration state, that is, fully solvated in explicit water, for a total system size of approximately 18,000 atoms. The power spectrum was calculated from the final 125 ps of a 1.6 ns simulation recorded at 2 fs resolution, to match the simulation length and sampling frequency employed in Ref [41]. This system has also been investigated with the AMBER classical force field and several machine learned potentials, including GEMS and, more recently, SO3LR [40, 39, 9]. All three MLPs produce spectra in qualitative agreement with each other. The water peaks at 1640 cm-1 and 3200-3600 cm-1 are reproduced reasonably accurately, especially considering that no nuclear quantum effects are taken into account (Figure 11). Furthermore, in the low frequency region, two peaks are identified in the power spectrum by GEMS, SO3LR and MACE-OFF24(M). The most obvious of these occurs at 60 cm-1 and corresponds to localized internal side chain fluctuations, while a peak at 220 cm-1 is due to specific configurations of water at the protein surface.
We note that, whilst a 125 ps trajectory is sufficient to qualitatively identify the spectral features, a much longer simulation of at least several nanoseconds would be required to fully converge the THz region of the spectrum and enable full quantitative comparison with experiment, as demonstrated for the GEMS model in Ref. [39]. We note that the power spectrum computed on the final 1.3 ns simulation (discarding the initial 300 ps to equilibration) does not yield a significantly different spectrum to that of the 125 ps simulation used in Figure 11, indicating that significantly more sampling would be required to resolve the THz lineshapes sufficiently to enable experimental comparison [87].